# From graph to Riesz continuity

### v2

@@ 4 Polarizations
Proof. Let us describe the plan of the proof. For an arbitrary polarization $\Pi$ we define a fiber bundle $\mathcal{T}=\mathcal{T}_{\Pi}$ over $X$ whose (local or global) sections are exactly (local or global) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}$ admits local sections if and only if $\Pi$ is tame. We will show that in this case the bundle $\mathcal{T}$ is locally trivial and has a contractible fiber. Such a bundle, over a paracompact base space $X$, admits global sections, and every section over a closed subset $Y\subset X$ can be extended to a global section [DD, Lemma 4]. Therefore, $\Pi$ admits global adapted trivializations, and every trivialization of $\mathcal{H}$ over $Y$ adapted to $\Pi$ can be extended to a global adapted trivialization.
Let $H_{X}$ be the trivial Hilbert bundle over $X$ with the fiber $H$. Consider the fiber bundle $\mathcal{U}(\mathcal{H},H_{X})$ associated with $\mathcal{H}$ whose sections are Hilbert bundle isomorphisms $\mathcal{H}\to H_{X}$ (that is, trivializations of $\mathcal{H})$. Every two such isomorphisms differ by an automorphism of $H_{X}$, which is defined by a continuous map $X\to{{}^{s}{\mathcal{U}}}(H)$. Therefore, $\mathcal{U}(\mathcal{H},H_{X})$ is a locally trivial principal bundle over $X$ with the structure group ${{}^{s}{\mathcal{U}}}(H)$. Its fiber over $x$ is ${{}^{s}{\mathcal{U}}}(\mathcal{H}_{x},H)$, and ${{}^{s}{\mathcal{U}}}(H)$ acts on the fiber by the left multiplication. A trivialization $\tau\colon\mathcal{H}\to H_{X}$ takes $\mathcal{U}(\mathcal{H},H_{X})$ to the trivial bundle $\mathcal{U}(H_{X},H_{X})=X\times{{}^{s}{\mathcal{U}}}(H)$.
The key ingredient of our proof is the bundle $\mathcal{T}\to X$ which we will now define. The total space $\mathcal{T}$ coincides with $\mathcal{U}(\mathcal{H},H_{X})$ as a set, and the projection to $X$ is the same, but we equip $\mathcal{T}$ with a finer topology determined by $\Pi$. Hence the identity map
(4.1) $\iota\colon\mathcal{T}\to\mathcal{U}(\mathcal{H},H_{X})$
will be continuous and, in particular, the projection $\mathcal{T}\to X$ will be continuous.
Suppose first that the polarization $\Pi$ is defined by a subbundle $\mathcal{E}$ of $\mathcal{H}$, that is, the atlas of $\Pi$ has only one chart $(X,\mathcal{E})$. Then $\mathcal{E}$ determines the map
(4.2) $\iota_{\mathcal{E}}\colon\mathcal{T}\to\operatorname{Gr}(H)\cong\mathcal{P}(H),\quad(x,u)\mapsto u(\mathcal{E}_{x})\;\;\text{for}\;\;x\in X\;\;\text{and}\;\;u\in\mathcal{U}(\mathcal{H}_{x},H).$
We equip $\mathcal{T}=\mathcal{T}_{\mathcal{E}}$ with the weakest topology for which both maps (4.1) and (4.2) are continuous.
If $\mathcal{E}$ is replaced by a compatible subbundle $\mathcal{F}$, then $\mathcal{T}_{\mathcal{F}}=\mathcal{T}_{\mathcal{E}}$. Indeed, in this case the difference $k_{x}=p_{x}-q_{x}$ of projections onto $\mathcal{F}_{x}$ and $\mathcal{E}_{x}$ is a norm continuous family of compact operators, so $\iota_{\mathcal{F}}$ differs from $\iota_{\mathcal{E}}$ by a continuous map
$\mathcal{U}(\mathcal{H},H_{X})\to\mathcal{K}(H),\quad(x,u)\mapsto uk_{x}u^{*}.$
Therefore, $\iota_{\mathcal{F}}$ and $\iota_{\mathcal{E}}$ induce the same refinement of the original topology on $\mathcal{U}(\mathcal{H},H_{X})$ and thus $\mathcal{T}_{\mathcal{E}}$ depends only on the equivalence class of compatible subbundles $\mathcal{E}$. This allows to define the bundle $\mathcal{T}_{\Pi}$ for an arbitrary polarization $\Pi$ of $\mathcal{H}$.
More precisely, let $\Pi$ be given by an atlas $(X_{i},\mathcal{E}_{i})$, $i\in I$. Then the restrictions of $\mathcal{T}_{\mathcal{E}_{i}}$ and $\mathcal{T}_{\mathcal{E}_{j}}$ to $X_{i}\cap X_{j}$ coincide. We equip $\mathcal{T}_{\Pi}$ with the topology defined by the requirement that its restriction to $X_{i}$ coincides with $\mathcal{T}_{\mathcal{E}_{i}}$ for every $i\in I$.
By the definition of $\mathcal{T}_{\Pi}$, its (local) sections are exactly (local) trivializations of $\mathcal{H}$ adapted to $\Pi$. In particular, $\mathcal{T}_{\Pi}$ admits local sections if $\Pi$ is tame. The following lemma provides a stronger result.
@@
If $\Pi$ is tame, then the bundle $\mathcal{T}_{\Pi}$ is locally trivial.
Proof. Since $\Pi$ is tame, each point $x_{0}\in X$ has a neighborhood $X^{\prime}$ such that some trivialization of $\mathcal{H}$ over $X^{\prime}$ is adapted to $\Pi$. We can suppose that $X^{\prime}$ is contained in some $X_{i}$; then this trivialization takes $\mathcal{E}_{i}$ to the subbundle $\mathcal{E}^{\prime}$ given by a continuous map $\mathcal{E}^{\prime}\colon X^{\prime}\to\operatorname{Gr}^{\star}(H)$. Let $g\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ be a map given by (2.1) and $\bar{g}_{L}=g_{L}^{-1}$. Then the composition of $\mathcal{E}^{\prime}$ with $\bar{g}\colon\operatorname{Gr}^{\star}\to\mathcal{U}(H)$ provides the automorphism $G$ of the trivial Hilbert bundle $H_{X^{\prime}}$ which takes $\mathcal{E}^{\prime}_{x}$ to the subspace $E$ of $H$ independent of $x$. We can then replace the original adapted trivialization by its composition with $G$. The new trivialization takes $\mathcal{E}_{i}$ to the subbundle $E_{X^{\prime}}$ of $H_{X^{\prime}}$ and thus provides a bundle isomorphism $\left.\mathcal{T}\right|_{X^{\prime}}\to\mathcal{T}^{\prime}$, where the bundle $\mathcal{T}^{\prime}$ is build (in the way described above) from the trivial bundle $H_{X^{\prime}}$ and its subbundle $E_{X^{\prime}}$. We will prove that the bundle $\mathcal{T}^{\prime}$ is trivial and find its fiber.
By the definition, the total space $\mathcal{T}^{\prime}$ is obtained from the product $X^{\prime}\times{{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that the map $\mathcal{T}^{\prime}\to\operatorname{Gr}^{\star}(H)$, $(x,u)\mapsto uE$ becomes continuous. This map does not actually depend on $x$. Therefore, $\mathcal{T}^{\prime}$ coincides with the trivial bundle $X^{\prime}\times T\to X^{\prime}$, where the fiber $T$ is obtained from ${{}^{s}{\mathcal{U}}}(H)$ by refining its topology so that
(4.3) $\iota_{E}\colon T\to\operatorname{Gr}^{\star}(H),\quad u\mapsto uE$
becomes continuous. In other words, $T$ has the topology induced by the inclusion
$T\hookrightarrow{{}^{s}{\mathcal{U}}}(H)\times\operatorname{Gr}^{\star}(H),\quad u\mapsto(u,uE).$
Hence the restriction of $\mathcal{T}_{\Pi}$ to a neighborhood of an arbitrary point $x_{0}$ is a trivial bundle with the fiber $T$. $\square$
Lemma 4.3.
The space $T$ defined above is contractible.
Proof. The fiber of $\iota_{E}$ over $E$ is the subspace of ${{}^{s}{\mathcal{U}}}(H)$ consisting of unitary operators preserving $E$. Denote this subspace by ${{}^{s}{\mathcal{U}}}_{E}$. We will show that (4.3) is a trivial bundle over $\operatorname{Gr}^{\star}$ and construct its trivialization $\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to T$.
Let $g\colon\operatorname{Gr}^{\star}(H)\to\mathcal{U}(H)$ be a continuous map such that $L=g_{L}E$ for all $L\in\operatorname{Gr}^{\star}(H)$, as in (2.1). We define $\varphi$ by the formula
$\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H)\;\to\;T,\quad\varphi(v,L)=g_{L}\cdot v.$
Since
(4.4) $(g_{L}v)E=g_{L}E=L\;\;\text{for every}\;\;v\in{{}^{s}{\mathcal{U}}}_{E},$
the composition $\iota_{E}\circ\varphi\colon{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}\to\operatorname{Gr}^{\star}$ is the projection of the product on the second factor; in particular, it is continuous. The composition of $\varphi$ with the identity map $T\to{{}^{s}{\mathcal{U}}}$ is given by $(v,L)\mapsto g_{L}v$ and is also continuous. Therefore, $\varphi$ itself is continuous.
$\psi\colon T\to{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}(H),\quad\psi(u)=\left(g_{uE}^{-1}\cdot u,\,uE\right).$
By the construction, $\psi\circ\varphi(v,L)=(v,L)$ for every $(v,L)\in{{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$. Conversely,
$\varphi\circ\psi(u)=\varphi\left(g_{uE}^{-1}\cdot u,\,uE\right)=g_{uE}\cdot g_{uE}^{-1}\cdot u=u\;\;\text{for every}\;\;u\in T.$
The domain of $\varphi$ is the product of two factors. The second factor $\operatorname{Gr}^{\star}$ is contractible, see [AS, proof of Lemma 3.6]. The first factor ${{}^{s}{\mathcal{U}}}_{E}={{}^{s}{\mathcal{U}}}(E)\times{{}^{s}{\mathcal{U}}}(E^{\bot})$ is the product of two copies of ${{}^{s}{\mathcal{U}}}$. By [DD, Lemma 3], ${{}^{s}{\mathcal{U}}}$ is contractible, so ${{}^{s}{\mathcal{U}}}_{E}$ is also contractible. It follows that the product ${{}^{s}{\mathcal{U}}}_{E}\times\operatorname{Gr}^{\star}$ is contractible. This completes the proof of the lemma. $\square$
End of the proof of Theorem 4.1.   Lemmas 4.2 and 4.3 together with the reasoning in the beginning of the proof show that every adapted trivialization of $\mathcal{H}$ over a closed subspace $Y\subset X$ can be extended to a global adapted trivialization. If $\tau_{0}$ and $\tau_{1}$ are trivializations of $\mathcal{H}$ adapted to $\Pi$, then $\tau_{0}\sqcup\tau_{1}$ is a trivialization of $\tilde{\mathcal{H}}$ over $X\times\left\{0,1\right\}$ adapted to $\tilde{\Pi}$, where $\tilde{\mathcal{H}}$ and $\tilde{\Pi}$ are the liftings of $\mathcal{H}$ and $\Pi$ to $X\times[0,1]$. The polarization $\tilde{\Pi}$ is tame, so $\tau_{0}\sqcup\tau_{1}$ can be extended from $X\times\left\{0,1\right\}$ to a trivialization of $\tilde{\mathcal{H}}$ adapted to $\tilde{\Pi}$. This provides a homotopy between $\tau_{0}$ and $\tau_{1}$ in the class of trivializations adapted to $\Pi$ and completes the proof of the theorem. $\square$
Finite covering dimension.   The proof of Theorem E in the next section is based on the following result.
Theorem 4.4.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a metrizable space $X$ of locally finite covering dimension. Then every polarization $\Pi$ of $\mathcal{H}$ is tame and thus satisfies all the conclusions of Theorem 4.1.
Proof. Every subspace $X^{\prime}$ of $X$ is metrizable and thus paracompact, and has locally finite covering dimension by [E, Theorem 7.3.1]. By [DD, Theorem 5], every separable Hilbert bundle over $X^{\prime}$ of infinite rank is locally trivial. The bundle $\mathcal{H}$ is trivial by [DD, Theorem 1] and has separable fiber, so it is separable in the sense of [DD], and so are its subbundles over $X^{\prime}$. Therefore, every infinite rank subbundle of $\mathcal{H}$ over $X^{\prime}$ is locally trivial.
Let $(X_{i},\mathcal{E}_{i})$ be an atlas of $\Pi$. By the reasoning above, both $\mathcal{E}_{i}$ and $\mathcal{E}_{i}^{\bot}$ are locally trivial. The direct sum of their local trivializations provides a local trivialization of $\mathcal{H}$ which takes $\mathcal{E}_{i}$ to a constant subspace and thus is adapted to $\mathcal{E}_{i}$. Therefore, $\Pi$ is tame. Since $X$ is paracompact, $\Pi$ satisfies the assumptions of Theorem 4.1. $\square$
Remark.
@@
For a general space $X$, such a result is no longer true. Indeed, Dixmier and Douady constructed a subbundle $\mathcal{E}$ of the trivial Hilbert bundle $\mathcal{H}$ over a compact space $X$, which is not locally trivial [DD, Theorem 6]. We recall their construction in Example 5.5 below. Every fiber of $\mathcal{E}$ has infinite rank and corank, so $\mathcal{E}$ determines a polarization of $\mathcal{H}$. However, $\mathcal{E}$ has no adapted trivializations, even locally.
@@ 5 Self-adjoint operators
Adapted trivializations.   We say that a (local or global) trivialization of $\mathcal{H}$ is adapted to a family $\mathcal{A}$ of regular operators $\mathcal{A}_{x}\in\mathcal{R}(\mathcal{H}_{x})$ if it takes $\mathcal{A}$ to a Riesz continuous family of operators in $H$. If operators $\mathcal{A}_{x}$ are bounded (say, they are projections), then an adapted trivialization takes $\mathcal{A}$ to a norm continuous family. So the definition from the previous section, of a trivialization adapted to a subbundle, agrees with the present definition.
Bounded operators.   Let $\mathcal{A}$ be a family of self-adjoint bounded operators acting on fibers of a Hilbert bundle $\mathcal{H}$. We want to know whether $\mathcal{A}$ admits a global adapted trivialization. Clearly, for that to be true, $\mathcal{A}$ has to be norm quasi-continuous (that is, admit local adapted trivializations). The following theorem shows that for essentially unitary operators this local condition is also sufficient. The author does not know whether this is true for arbitrary bounded self-adjoint operators.
Theorem 5.1.
Let $\mathcal{H}$ be a locally trivial Hilbert bundle over a paracompact space $X$ and $\mathcal{A}$ be a family of self-adjoint essentially unitary operators. Suppose that $\mathcal{A}$ is norm quasi-continuous. Then there exists a global trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$. Moreover, every adapted trivialization of $\mathcal{H}$ over a closed subset $Y$ of $X$ can be extended to a global adapted trivialization. In particular, every two such trivializations are homotopic (in the space of trivializations adapted to $\mathcal{A}$).
Proof. Let $X^{+}$, resp. $X^{-}$ be the subset of points $x\in X$ for which $\mathcal{A}_{x}$ is essentially positive, resp. negative, and let $X^{\star}=X\setminus(X^{+}\cup X^{-})$. Since $\mathcal{A}$ is norm quasi-continuous, each of the three subspaces $X^{+}$, $X^{-}$, and $X^{\star}$ is open and thus closed in $X$, so it is enough to prove the theorem for each of these subspaces separately.
For $x\in X^{+}$, the operator $\mathcal{A}_{x}$ is essentially positive and essentially unitary, so $\mathcal{A}_{x}-1$ is compact. Over $X^{+}$ the family $x\mapsto\mathcal{A}_{x}-1$ is a norm quasi-continuous family of compact operators, so both $\mathcal{A}-1$ and $\mathcal{A}$ are norm continuous in every trivialization of $\mathcal{H}$ over $X^{+}$. The similar reasoning shows that $\mathcal{A}$ is norm continuous in every trivialization of $\mathcal{H}$ over $X^{-}$. Trivializations of $\mathcal{H}$ over $X^{+}\cup X^{-}$ are identified with maps $X^{+}\cup X^{-}\to{{}^{s}{\mathcal{U}}}(H)$, so every two trivializations are homotopic.
@@
(5.1) $p_{\lambda}(x)=\operatorname{\mathbb{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})\;\;\text{for}\;\;x\in X_{\lambda}=\left\{x\in X\,|\,\lambda\in\operatorname{Res}(\mathcal{A}_{x})\right\}.$
The sets $X_{\lambda}$ with $\lambda$ running the interval $(-1,1)$ form an open covering of $X$. A local trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$ takes $p_{\lambda}$ to a norm continuous family over $X_{\lambda}$. For $\lambda<\mu$ the difference
(5.2) $p_{\lambda}(x)-p_{\mu}(x)=\operatorname{\mathbb{1}}_{[\lambda,\mu)}(\mathcal{A}_{x})=\operatorname{\mathbb{1}}_{[\lambda,\mu]}(\mathcal{A}_{x})$
is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$. Hence the collection $(X_{\lambda},p_{\lambda})$, $\lambda\in(-1,1)$ is an atlas of a tame polarization $\Pi_{\mathcal{A}}$ of $\mathcal{H}$ over $X$. The difference $k_{\lambda}=\mathcal{A}-(2p_{\lambda}-1)$ is a norm continuous family of compact operators over $X_{\lambda}$ in a local trivialization adapted to $\mathcal{A}$, so it is norm continuous in every trivialization of $\mathcal{H}$ over $X_{\lambda}$. A trivialization of $\mathcal{H}$ over $X_{\lambda}$ is adapted to $\mathcal{A}=k_{\lambda}+(2p_{\lambda}-1)$ if and only if it is adapted to $p_{\lambda}$. Hence a trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4.1. $\square$
Remark.
In fact, the main part of Theorem 5.1, about operators which are neither essentially positive nor essentially negative, is equivalent to Theorem 4.1. Indeed, let $(X_{i},p_{i})$ be an atlas of a polarization $\Pi$. We associate with it the family $\alpha$ of self-adjoint essentially unitary operators $\alpha(x)=\sum_{i}\rho_{i}(x)r_{i}(x)$, where $(\rho_{i})$ is a partition of unity subordinated to the covering $(X_{i})$ and $r_{i}=2p_{i}-1$. Then $\alpha-r_{i}=\sum_{j}\rho_{j}(r_{j}-r_{i})$ is a norm continuous family of compact operators over $X_{i}$. Therefore, a (local or global) trivialization of $\mathcal{H}$ is adapted to $\Pi$ if and only if it is adapted to $\alpha$.
Proof of Theorem D.   The bounded transform takes self-adjoint operators with compact resolvents to self-adjoint essentially unitary operators and Riesz (quasi-)continuous families to norm (quasi-)continuous families. Therefore, Theorem D follows immediately from Theorem 5.1.
Spectral projections.   For a family $\mathcal{A}_{x}$, $x\in X$ of self-adjoint operators and $\lambda\in\mathbb{R}$ we consider the family $p_{\lambda}$ of spectral projections over $X_{\lambda}\subset X$ defined by (5.1).
@@
If $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ is graph continuous, then $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is strong continuous for every $\lambda\in\mathbb{R}$.

### v3

@@ 6 Proof of Theorem 5
Let $X$ be a paracompact space of finite covering dimension. Then every continuous map $p\colon X\to{{}^{s}{\mathcal{P}}}^{\star}(H)$ can be lifted to a map $u\colon X\to{{}^{s}{\mathcal{U}}}(H)$. In other words, every subbundle of $H_{X}$ with infinite rank and corank is trivial.
Proof. Let $S(H)$ be the unit sphere in $H$. For $q\in\mathcal{P}(H)$ and $F=\operatorname{Ran}q$ we denote the unit sphere in $F$ by $S(q)=S(F)$.
The map $p\colon X\to{{}^{s}{\mathcal{P}}}^{\star}(H)$ determines a bundle projection $H_{X}\to H_{X}$ of the trivial bundle $H_{X}$, which we also denote by $p$. Let $\mathcal{E}$ be the range of $p$ (it is a subbundle of $H_{X}$) and
$S(p)=S(\mathcal{E}):=(X\times S(H))\cap\mathcal{E}=\left\{(x,e)\mid x\in X,\;e\in S(\mathcal{E}_{x})\right\}$
be the bundle of unit spheres in $\mathcal{E}$.
@@
Let $\eta$ be a continuous section of $\mathcal{E}\to X$ and $\varepsilon>0$. Then there is a continuous section $\xi$ of $S(\mathcal{E})\to X$ such that the distance from $\eta(x)$ to the linear span $\left\langle\xi(x)\right\rangle$ of $\xi(x)$ is less than $\varepsilon$ for every $x\in X$.
Proof. For a non-zero vector $v\in H$, we will denote by $\mathbf{s}(v)=\lVert v\rVert^{-1}\cdot v\in S(H)$ the corresponding unit vector.
Consider the family $\mathcal{S}=\left\{S(q)\mid q\in\mathcal{P}^{\star}(H)\right\}$ of closed subsets of $H$. Let us check that the composition $X\to{{}^{s}{\mathcal{P}}}^{\star}\to\mathcal{S}$ taking $x\in X$ to $S(p_{x})\in\mathcal{S}$ satisfies all four assumptions of Michael’s Selection Theorem [M, Theorem 1.2].
(1) $X$ is paracompact and has finite covering dimension.
(2) Every element of the family $\mathcal{S}$ is contractible. Indeed, the unit spheres $S(q)$ are contractible by Kakutani’s theorem [Kak, Theorem 3].
(3) Let us show that the family $\mathcal{S}$ is equi-$LC^{n}$ in terms of [M] for every $n$. In fact, the stronger condition holds: for every open ball $V$ of radius <1<1 with the center in $v_{0}\in S(H)$ and for every $q\in\mathcal{P}^{\star}$, the intersection $V\cap S(q)$ is either empty or contractible. Indeed, if $V\cap S(q)$ is non-empty, then it coincides with $V^{\prime}\cap S(q)$, where $V^{\prime}$ is an open ball of the same or smaller radius with the center in $v^{\prime}_{0}=\mathbf{s}(qv_{0})\in S(q)$. The formula $(v,t)\mapsto\mathbf{s}((1-t)v+tv^{\prime}_{0})$ provides a contraction of $V^{\prime}\cap S(q)$. Therefore, $\mathcal{S}$ is equi-$LC^{n}$.
(4) Let us show that the map ${{}^{s}{\mathcal{P}}}^{\star}(H)\to\mathcal{S}$, $q\mapsto S(q)$ is lower semi-continuous, that is, for every open $V\subset H$ the set $U=\left\{q\in\mathcal{P}^{\star}\mid S(q)\cap V\neq\emptyset\right\}$ is open in $^{s}{\mathcal{P}}$$^{\star}$. Let $q_{0}\in U$ and $v_{0}\in S(q_{0})\cap V$. Then $W=\left\{q\in\mathcal{P}^{\star}\mid\lVert(q-q_{0})v_{0}\rVert<1\right\}$ is a neighborhood of $q_{0}$ in $^{s}{\mathcal{P}}$$^{\star}$ and the map $f\colon W\to H$, $f(q)=\mathbf{s}(qv_{0})\in S(q)$ is continuous and has the prescribed value $v_{0}$ at $q_{0}$. Thus the inverse image $f^{-1}(V)\subset W$ is open in $^{s}{\mathcal{P}}$$^{\star}$, contains $q_{0}$, and is contained in $U$. Therefore, $U$ is open in $^{s}{\mathcal{P}}$$^{\star}$.
We see that the map $X\to\mathcal{S}$, $x\mapsto S(p_{x})\in\mathcal{S}$ satisfies assumptions of [M, Theorem 1.2] and thus every section of $S(\mathcal{E})$ over a closed subset $Y\subset X$ can be extended to a global section of $S(\mathcal{E})$.
Let $Y=\left\{x\in X\mid\lVert\eta(x)\rVert\geqslant\varepsilon\right\}$. Then the vector field $\xi\colon Y\to S(H)$, $\xi(x)=\mathbf{s}(\eta(x))$ is a section of $S(\mathcal{E})$ over $Y$. Let $\xi$ be an extension of this section to the whole $X$. By the construction, the distance from $\eta(x)$ to $\left\langle\xi(x)\right\rangle$ is zero for $x\in Y$ and less than $\varepsilon$ for $x\notin Y$, so $\xi$ is a required section. $\square$
Continuation of the proof of Theorem 6.1.  We aim to construct continuous maps $\xi_{n}\colon X\to S(H)$, $n\in\mathbb{N}$ such that $\xi_{1}(x),\xi_{2}(x),\ldots$ is an orthonormal basis of $\mathcal{E}_{x}$ for every $x\in X$. To this end, fix a countable dense subset $\bar{\eta}_{1},\bar{\eta}_{2},\ldots$ of $H$. Then $p\bar{\eta}_{n}$ are sections of $\mathcal{E}\to X$ and the set $\left\{p(x)\bar{\eta}_{n}\right\}_{n\in\mathbb{N}}$ is dense in $\mathcal{E}_{x}$ for every $x$. We will now construct bundle projections $p_{1}\geqslant p_{2}\geqslant\ldots$ of $H_{X}$ and sections $\xi_{n}$ of $S(p_{n})$ by induction as follows. We start with $p_{1}=p$. Suppose that $p_{n}$ is already constructed. Applying the lemma above to $\varepsilon_{n}=1/n$, $p_{n}$, and $\eta_{n}=p_{n}\bar{\eta}_{n}$, we obtain a continuous section $\xi_{n}\colon X\to S(p_{n})$ such that the distance from $\eta_{n}(x)$ to $\left\langle\xi_{n}(x)\right\rangle$ is less than $\varepsilon_{n}$ for every $x\in X$. Let $q_{n}(x)$ be the orthogonal projection onto $\left\langle\xi_{n}(x)\right\rangle$. Then the map $q_{n}\colon X\to\mathcal{P}(H)$ is norm continuous, and we proceed with the next bundle projection $p_{n+1}=p_{n}-q_{n}$. In such a way, we obtain continuous sections $\xi_{1},\xi_{2},\ldots$ of $S(\mathcal{E})$ such that the vectors $\xi_{n}(x)$ are pairwise orthogonal and their finite linear combinations are dense in $\mathcal{E}_{x}$ for every $x\in X$.
The same reasoning applied to 1p1-p instead of $p$ provides continuous sections $\xi^{\prime}_{1},\xi^{\prime}_{2},\ldots$ of $S(H_{X}\ominus\mathcal{E})$ such that for every $x$ the vectors $\xi^{\prime}_{n}(x)$ form an orthonormal basis of $H\ominus\mathcal{E}_{x}$.
Fix an orthonormal basis $e_{1},e_{2},\ldots$ of $E$ and an orthonormal basis $e^{\prime}_{1},e^{\prime}_{2},\ldots$ of $H\ominus E$. Let $u(x)\colon H\to H$ be the isometry given by the rule $u(x)e_{n}=\xi_{n}(x)$ and $u(x)e^{\prime}_{n}=\xi^{\prime}_{n}(x)$. Then the range of $u(x)$ is dense in $H$, so $u(x)$ is unitary. By the construction, $x\mapsto u(x)$ is strongly continuous and $u(x)E=\mathcal{E}_{x}$, so the map $u\colon X\to{{}^{s}{\mathcal{U}}}(H)$ lifts $p\colon X\to{{}^{s}{\mathcal{P}}}(H)$.
The trivialization of the trivial Hilbert bundle $H_{X}$ given by $u^{-1}$ takes $\mathcal{E}$ to the constant subbundle $E_{X}$ of $H_{X}$ and thus is adapted to $\mathcal{E}$. $\square$
Proof of Theorem 5.  The structure group $^{s}{\mathcal{U}}$} of a locally trivial Hilbert bundle $\mathcal{H}$ is contractible [DD, Lemma 3]. Since the base space $X$ is paracompact, $\mathcal{H}$ is trivial.
Let $(X_{i},\mathcal{E}_{i})$ be an atlas of a polarization $\Pi$. Let $x\in X_{i}$ be an arbitrary point. By the assumption of the theorem, $x$ has a neighborhood $X^{\prime}$ of finite covering dimension. Since $X$ is normal, there is a closed neighborhood $Y$ of $x$ which is contained in $X^{\prime}\cap X_{i}$. Such an $Y$ is a closed subset of a paracompact space $X$, so it is also paracompact; it is a closed subset of $X^{\prime}$, so $\dim Y\leqslant\dim X^{\prime}<\infty$. By Theorem 6.1, the restriction of $\mathcal{E}_{i}$ to $Y$ is trivial. Since $x$ was chosen arbitrarily, each $\mathcal{E}_{i}$ is tame. Therefore, $\Pi$ is also tame, so it satisfies the assumptions of Theorem 4 and thus admits a global adapted trivialization. $\square$
Remark 6.3.
@@
For a general space $X$, such a result is no longer true. Indeed, Dixmier and Douady constructed a subbundle $\mathcal{E}$ of the trivial Hilbert bundle $\mathcal{H}$ over a compact space $X$, which is not locally trivial. We recall their construction in Example 7.5 below. Every fiber of $\mathcal{E}$ has infinite rank and corank, so $\mathcal{E}$ determines a polarization of $\mathcal{H}$. However, $\mathcal{E}$ has no adapted trivializations, even locally.
@@ 7 Self-adjoint operators
Proof of Theorem 6.  A regular operator has compact resolvents if and only if its bounded transform is essentially unitary. A trivialization is adapted to a family $\mathcal{A}$ of regular operators if and only if it is adapted to the bounded transform $\bm{\chi}(\mathcal{A})$. Therefore, we only need to prove Theorem 6 for essentially unitary operators $\mathcal{A}_{x}$; the case of operators with compact resolvents will immediately follow.
Let $X^{+}$, resp. $X^{-}$ be the subset of points $x\in X$ for which $\mathcal{A}_{x}$ is essentially positive, resp. negative, and let $X^{\star}=X\setminus(X^{+}\cup X^{-})$. In a local adapted trivialization $\mathcal{A}$ is norm continuous, so each of the three subspaces $^{+}, XX^{-}$, and $X^{\star}$ is open in $X$. Therefore, $X$ is the disjoint union of $^{+}, XX^{-}$, and $X^{\star}$, so it is enough to prove the theorem for each of these subspaces separately.
For $x\in X^{+}$, the operator $\mathcal{A}_{x}$ is essentially positive and essentially unitary, so $\mathcal{A}_{x}-1$ is compact. Over $X^{+}$ the family $x\mapsto\mathcal{A}_{x}-1$ is a norm continuous family of compact operators, so both $\mathcal{A}-1$ and $\mathcal{A}$ are norm continuous in every trivialization of $\mathcal{H}$ over $X^{+}$. The similar reasoning shows that $\mathcal{A}$ is norm continuous in every trivialization of $\mathcal{H}$ over $X^{-}$. Trivializations of $\mathcal{H}$ over $X^{+}\cup X^{-}$ are identified with maps $X^{+}\cup X^{-}\to{{}^{s}{\mathcal{U}}}(H)$, so every two trivializations are homotopic.
@@
(7.1) $p_{\lambda}(x)=\operatorname{\mathds{1}}_{[\lambda,+\infty)}(\mathcal{A}_{x})\;\;\text{for}\;\;x\in X_{\lambda}=\left\{x\in X\mid\lambda\in\operatorname{Res}(\mathcal{A}_{x})\right\}.$
The sets $X_{\lambda}$ with $\lambda$ running the interval (1,1)(-1,1) form an open covering of $X$. A local trivialization of $\mathcal{H}$ adapted to $\mathcal{A}$ takes $p_{\lambda}$ to a norm continuous family over $X_{\lambda}$. For $\lambda<\mu$ the difference
(7.2) $p_{\lambda}(x)-p_{\mu}(x)=\operatorname{\mathds{1}}_{[\lambda,\mu)}(\mathcal{A}_{x})=\operatorname{\mathds{1}}_{[\lambda,\mu]}(\mathcal{A}_{x})$
is a norm continuous family of finite rank operators over $X_{\lambda}\cap X_{\mu}$. Hence the collection $(X_{\lambda},p_{\lambda})$, $\lambda\in(-1,1)$ is an atlas of a tame polarization $\Pi_{\mathcal{A}}$ of $\mathcal{H}$ over $X$. The difference $k_{\lambda}=\mathcal{A}-(2p_{\lambda}-1)$ is a norm continuous family of compact operators over $X_{\lambda}$ in a local trivialization adapted to $\mathcal{A}$, so it is norm continuous in every trivialization of $\mathcal{H}$ over $X_{\lambda}$. A trivialization of $\mathcal{H}$ over $X_{\lambda}$ is adapted to $\mathcal{A}=k_{\lambda}+(2p_{\lambda}-1)$ if and only if it is adapted to $p_{\lambda}$. Hence a trivialization is adapted to $\mathcal{A}$ if and only if it is adapted to $\Pi_{\mathcal{A}}$. It remains to apply Theorem 4. $\square$
Spectral projections.  For a family $\mathcal{A}_{x}$, $x\in X$ of self-adjoint operators and $\lambda\in\mathbb{R}$ we consider the family $p_{\lambda}$ of spectral projections over the subset $X_{\lambda}\subset X$ defined by (7.1).
@@
If $\mathcal{A}\colon X\to\mathcal{R}^{\mathrm{sa}}(H)$ is graph continuous, then $p_{\lambda}\colon X_{\lambda}\to\mathcal{P}(H)$ is strongly continuous for every $\lambda\in\mathbb{R}$.